OpenStudy (yanni):

Graph the line with slope -1/3 passing through the point (-3,5).

OpenStudy (anonymous):

use y=mx+b your slope -1/3 is m so y = -1/3x + b now plug in your x and y to solve for b (:

OpenStudy (yanni):

thx

OpenStudy (anonymous):

or use y-y1=m(x-x1) and solve for y. :)